In the global system, a point is defined by P = [-5; 8; -4]. A camera is positioned in the global system at [10; two; 3] and is focused on the point [0; 1; -3], with the UP vector in the [0; 1; 1]. How much is this point worth in the camera system?


1
Expert's answer
2021-05-12T18:03:20-0400

Given,

Co-ordinates of the points = [-5, 8, -4]

Co-ordinates of the camera = [10,2, 3]

Co-ordinates of the focus =[0,1,-3]

Position vector of the object = [0,1,1]

focal length "(f)=10\hat{i}+\hat{i}+6\hat{j}"

"|f|=\sqrt{137}"

"v= {15}\hat{i}-6\hat{j}+7\hat{k}"

"|v|=\sqrt{225+36+49}"

"=\sqrt{305}"

"u=\frac{fv}{v-f}"


"=\frac{\sqrt{305}\times \sqrt{137}}{\sqrt{305}-\sqrt{137}}"

Hence, m= -v/u

"=\frac{\sqrt{305}-\sqrt{137}}{\sqrt{137}}"